Uver is a ride-sharing service that has become very popular in large cities. In one of these large cities, Uver is trying a new model: selling a bundle of 10 rides per week (instead of charging for each ride). Let ? stand for the number of bundles, and suppose the weekly cost of providing a bundle of 10 rides is given by the following function: ?(?)=?33−10?2+200?+500

Assume all Uver drivers have the same cost function.

The market for Uver rides in this city is competitive. Suppose there are 174 Uver drivers in the city. If the price of a bundle of rides is $181, the number of bundles of rides that will be provided is bundles.
Answer #1

Ans) 1 bundle has 10 rides per week. Y stands for number of bundles.

Cost per bundle, c(y) = y33-10y^2+200y+500. Price per bundle or Marginal Revenue = 181.

In a perfectly competative market : MR = MC. -------------equation 1

Marginal cost or MC = d/dx (y33-10y^2+200y+500) ---------differentiation of total cost

MC = 33-20y+200

181 = 33-20y+200 ------------------as per equation 1

20 y = 233-181

y =52/20 = 2. 6 rounded of to 3 bundles or rides